Status Report

ISS On-Orbit Status 10 Jan 2003

By SpaceRef Editor
January 10, 2003
Filed under , ,
ISS On-Orbit Status 10 Jan 2003

All ISS systems continue to function nominally, except as noted previously or below.

At wake-up (1:00 am EST), crewmembers were congratulated on their great job on yesterday’s SSRMS/robotarm activities.  MCC-H and CSA (Canadian Space Agency) received a lot of excellent data and “wonderful” views of the external surveys.

FE-1 Nikolai Budarin started his day by laboring through a third Russian MBI-8 Profilaktika (preventive health maintenance) fitness session, today on the U.S. TVIS treadmill, with blood analysis.  The collected data were then transferred by Budarin from cardiocassette-2000 storage for downlink via OCA.  [This test is similar to the ÌÎ-3 test performed in TVIS passive mode with arbitrarily selected speeds (within a specified range). The difference from the nominal test is the use of the TEEM-100M gas analyzer during operations, blood lactate determination and subjective load assessment.  Don Pettit again provided CMO (crew medical officer) assistance, drawing Nikolai’s blood and taking digital imagery of him in gear on the TVIS.  The imagery file was then copied to the “Pulse” data memory card from the Pulse Kit for subsequent downlinking.]

Budarin also completed his daily checkout of the BIO-5 Rasteniya-2/Lada-2 (“Plants-2”) payload which studies plant growth and development under spaceflight conditions.

Later, Nikolai was scheduled to work on replacement of the Russian “Wiener” power laptop with an ISS Wiener computer.  [The task included test activation and reconfiguration of the new laptop and its software, and functionality tests of SSR (SmartSwitch router) control of the onboard Ethernet crew support network in the Russian segment, as well as of the network itself using the ISS Wiener.  The new computer was to be left on to monitor SSR.]

The crew continued preparations for next week’s Stage 11A EVA (1/15), including preparation of the Airlock (A/L) equipment lock, installation of Metox (metal oxide) regenerative CO2 removal canisters, preparation of EVA tools, and review of spacewalk timeline, procedures and checklist updates uplinked by MCC-H.  This was followed by an EVA conference with ground specialists.  Timeline challenge: Crew preparation time over the coming weekend should be maximized, with sufficient time allotted for Nikolai Budarin to familiarize himself with his new role as IV (intravehicular crewmember), e.g., assisting in EMU donning/doffing.  [The EVA by Bowersox and Pettit is to begin at about 7:30am EST, to last 6.5 hours (egress time impacts radiation exposure).  The SSRMS will not be used but utilized as camera platform from parked position.  Prior to the EVA, the MT (mobile transporter) will be moved from WS7 (work station 7) at the end of the P1 truss to WS4, allowing sufficient time for crew assist if required (to install MT extension cord).  Planned tasks are:  (1) configure SFU (squib firing unit) on P1 to launch condition; (2) release 10 launch restraint locks (5 zenith, 5 nadir) on the P1 radiator beam; (3) inspect radiator beam and monitor its deployment (with manual intervention if required); (4) reconfigure SFU; (5) clean Node nadir ACBM (active common berthing module); (7) preposition an APFR (articulating portable foot restraint) on the CETA (crew equipment transfer aid) cart, and (8) perform get-ahead tasks as open time permits.  It will be the 50th EVA for ISS assembly, with 25 from Shuttle and 25 from station (16 of these from the Joint Airlock, but it will be only the second with no Orbiter present).]

Budarin conducted another inspection of the Russian BRPK-2 condensate water separator, as BRPK-1 of the SRVK-2M condensate water processor continues failed.  SRVK is operating on one lane only (SRVK-2).  [Tomorrow, MCC-Moscow plans to have the CFU (condensate feed unit) removed and replaced with a spare, but attempts at repairing the basic problem of the water processing system, whose separator unit is getting “overwhelmed” with water, have not been successful to date.  RSC-Energia is considering a new, more capable CFU for future delivery to the ISS.]

Meanwhile, as a consequence, U.S. condensate water cannot be processed by the Russian SRVK system (an issue that will likely remain open for a long time).  CDR Ken Bowersox took a CWC (contingency water container) inventory today, and Budarin performed CWC servicing.  [The Lab condensate tank is approaching its upper limit fill capacity of 100 lbs (currently at >90 lb).  The plan of venting the excess water overboard, originally intended for 1/12 (Saturday-Sunday), has been changed.  Instead, the condensate will be offloadedinto CWCs  #1022 and #1004, whichalready contain approximately 20 liters each. This offload will give the U.S. condensate tank room enough to collect condensate for about two weeks.  After that, plans are to dump the unprocessed contents of the tank and two CWCs overboard while in XPOP, in retrograde direction (which slows the ejected mass down).]

FE-2/SO Don Pettit continued his daily monitoring of CO2 partial pressure levels in the SM and Lab (two readings) with the CDMK (carbon dioxide monitoring kit).  Both the SM gas analyzer and the U.S. MCA (major constituents analyzer) have been given suspect values.

The two U.S. crewmembers performed their regular daily physical exercise program on TVIS, RED and CEVIS.

Yesterday’s R&R of RPCM (remote power controller module) N14B in the Node was completed as planned, but one of the RPC failed open after two tries.  The unit was powering one string of MT (mobile transporter) heaters, which are now operating on the remaining string.

After yesterday’s replacement of its VN vacuum pump, the Vozdukh CO2 removal system is working normally.  Consequently, the U.S. CDRA (carbon dioxide removal assembly) was deactivated yesterday, and it remained for the crew today to disconnect the ITCS LTL (internal thermal control system low temperature loop) supply line.  [The LTL jumpermust be disconnected within24 hours after CDRAdeactivation.]

Bowersox downloaded stored exercise files data from the TVIS treadmill, RED exerciser and CEVIS cycle ergometer to the MEC (medical equipment computer), a regular weekly task.  Later, Pettit transferred CEVIS data from the MEC to a PCMCIA (personal computer memory card international association) card for return to Earth.

TsUP/MCC-M reported that the newly installed 800A storage battery #5 in the Service Module (SM) is fully operational.

TsUP also reported that a test of the SM BITS onboard telemetry measurement system yesterday brought up a warning message of the Elektron system’s FDIR (failure detection, isolation, and recovery) unit.  This was apparently unrelated to the Elektron hardware but caused by some transient software glitch.

U.S. CMG-4 (control moment gyroscope #4) last night exhibited an increase in spin motor current from 0.47 to 0.53 amps, then stabilized at that level.  [The increase, which was coincident with heater activation, has been seen before prior to XPOP attitude.  Station attitude is now back at LVLH (local vertical/local horizontal), with somewhat higher bearing lubricant viscosity and commensurately elevated spin motor current.  Nevertheless, the CMG is being watched carefully, and if current reaches 0.57 amps, further steps will be taken.  Spin-down of the gyro is required at a current level of 0.6 amps, which would limit operations to two gyros.  The station can be controlled with two CMGs until Mission 12A (May ’03), and a spare CMG is manifested to be delivered prior to that, by ULF-1 in March.]

The RED (resistive exercise device) continues to be of concern, and further IFM (in-flight maintenance) is being discussed.  As encountered previously, one canister is exhibiting clicking noises, and both cans have a problem with pull cord retraction.

Three new U.S. HDDs (hard disk drives) have safely arrived in Moscow, to be carried  to Baikonur on 1/14, in time to be loaded on Progress 10P (scheduled launch: 2/2).  [Testing of the HDDs, which contain software serving as crew interface for the next station software upgrade next month, has been successfully completed, and processing of final safety certification paperwork is underway.]

Preparation of the replacement parts for the MSG (microgravity science glovebox) has proceeded on a tight schedule at ESA and its contractor, Astrium.  Of the two returned parts, the ESEM3 (exchangeable standard electronic module 3) did not require any changes, but the PDC (power distribution controller) was modified, along with a second PDC.  Some testing and verification work is still required, but the parts are expected to be in Baikonur in time before Progress 10P is closed out on 1/25 for irreversible liquid loading ops.  [PDC electrical functional test was completed yesterday, thermal cycling testing is to be done over this weekend, vibration & post-vibration testing is planned for 1/13, stand-alone environmental/electromagnetic interference testing at Astrium follows at 1/14-16, with subsequent hand-carry to MSFC/Huntsville for integrated engineering testing & recertification in the ground unit until 1/21, then hand-carry back to ESA and Russia on 1/22-1/24.  The question whether both PDCs should be tested before delivery to MSFC, or only one at a time (to possibly reduce time risk) is under discussion.]

At 1:18pm EST, the crew conducted an interactive educational TV event with middle school students from around Nevada, assembled at KLVX TV studio in Las Vegas, NV.  The program was taped and will later be aired by KLVX TV and internal broadcast to rural, urban and at-risk Nevada students.

Today’s targets for the CEO (crew earth observations program) were Baghdad, Iraq (nadir pass; ESC [electronic still camera]), Kuwait City, Kuwait (nadir and a touch right; ESC), Bamako, Mali (nadir pass; ESC. Fires in the Sahel continue making the news.  Crew was to continue to document [views left and right, especially obliques]), Angolan Biomass Burning (high pressure advancing from the west. Shoot obliques for any aerosols.  Views mainly right), Industrialized SE Africa (high pressure advancing from the west. Shooting obliques for any aerosols.  Views mainly right), Dallas, Texas (nadir pass; ESC), and Panama Canal (DYNAMIC EVENT:  Unusually cloud free.  Looking right of track ~2.5 degrees). CEO images can be viewed at the website

U.S. and Russian Segment Status (as of 1:10pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On (24 Amp mode).  Vozdukh CO2 scrubber is On.  U.S. CDRA CO2 scrubber is Off. TCCS is operational.  BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, bed #2 in Purify mode.
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 24.5; ppO2 (mmHg) — 159.5; ppCO2 (mmHg) — 0.9 (suspect).
  • SM Transfer Compartment:  Pressure (mmHg) — 751; temperature (deg C) — 19.1.
  • FGB Cabin:  Pressure (mmHg) — 752; temperature (deg C) — 21.0.
  • Node:  Pressure (mmHg) — 737.53; temperature (deg C) — 22.0 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 739.56; temperature (deg C) — 20.9; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
  • Joint Airlock (Equip. Lock):  Pressure (mmHg) — 739.66, temperature (deg C) — 24.9; shell heater temp (deg C) — 24.4, ppO2 (mmHg) — 158.5; ppCO2 (mmHg) — 3.6 (suspect).
  • PMA-1:  Shell heater temp (deg C) — 22.9
  • PMA-2:  Shell heater temp (deg C) — 17.4
  • (n/a = data not available)

Propulsion System (PS): 

  • Total propellant load available (SM + FGB + Progress) — 3689 kg (8133 lb) [as of 1/9/03].

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  Beta Gimbal Assembly (BGA) 2B and BGA 4B both in Dual-angle mode (directed position).
  • SM batteries (3am):  Battery #7 is off line (failed); battery #6 is in “Cycle” mode; all other batteries (6) are in “Partial Charge” mode.
  • FGB batteries (3am):  Batteries #3 is offline (failed); battery #4 is in “Cycle” mode; all other batteries (4) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 is in Standby; PCU-2 is in Standby.

Thermal Control Systems:

  • Air conditioner SKV-1 is Off, SKV-2 is Off.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup (new patches loaded on both).
  • EXT-2 is On (primary), EXT-1 is off.
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is On (primary); PL-2 MDM is off (diagnostic
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational. Lane 1 is down (as of 11/14).
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Source:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector — U.S. SIGI-1
  • Attitude — U.S. SIGI-1
  • Angular rates — U.S. RGA-1

Flight Attitude:

  • LVLH TEA (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -9.1 deg, roll: 0 deg]).
  • Solar Beta Angle:  -0.59 (magnitude increasing)

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operating.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally.
  • Ku-band is operating nominally.
  • Audio subsystem operating nominally.
  • Video subsystem operating nominally.
  • MCOR (medium-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings.  MT: at WS7, with KA power.  POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is Off; Cupola RWS is Off.

ISS Orbit (as of this morning, 7:21am EST [= epoch]):

  • Mean altitude — 391.0 km
  • Apogee — 394.5 km
  • Perigee — 387.5 km
  • Period — 92.38 min.
  • Inclination (to Equator) — 51.63 deg
  • Eccentricity — 0.0005199
  • Orbits per 24-hr. day — 15.59
  • Solar Beta Angle — -0.59 deg (magnitude increasing)
  • Altitude loss — 180 m (mean) in last 24 hours
  • Revolutions since FGB/Zarya launch (Nov. ’98) — 23640
  • For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.